# Learning to Translate Noise for Robust Image Denoising

### v1

@@ Abstract
Deep learning-based image denoising techniques often struggle with poor generalization performance to out-of-distribution real-world noise. To tackle this challenge, we propose a novel noise translation framework that performs denoising on an image with translated noise rather than directly denoising an original noisy image. Specifically, our approach translates complex, unknown real-world noise into Gaussian noise, which is spatially uncorrelated and independent of image content, through a noise translation network. The translated noisy images are then processed by an image denoising network pretrained to effectively remove Gaussian noise, enabling robust and consistent denoising performance. We also design well-motivated loss functions and architectures for the noise translation network by leveraging the mathematical properties of Gaussian noise. Experimental results demonstrate that the proposed method substantially improves robustness and generalizability, outperforming state-of-the-art methods across diverse benchmarks. Visualized denoising results and the source code are available on our project page.
@@ 1 Introduction
Image denoising aims to restore the pure signal from noisy images and serves as a critical preprocessing step to improve the visual quality of input images, extending the applicability of various downstream tasks. Recent advances in deep learning have significantly improved the performance of image denoising models [39, 40, 9, 35, 37, 36, 6, 38]. A common assumption in early approaches was that camera noise could be modeled as Gaussian noise [23, 39, 40], which simplified the process of generating noisy-clean image pairs by adding synthetic Gaussian noise. This allowed for the creation of large datasets that could be used to train denoising models in a supervised manner, playing a crucial role in advancing the development of denoising models.
Although these models trained on synthetic dataset perform well under controlled environments, they often struggle to generalize to real-world scenarios due to the fundamental differences between synthetic and real noise distributions [9]. In response, researchers have collected clean-noisy image pairs from real images [1, 31, 33] to address realistic noises, but models trained on this data still tend to overfit to the specific noise-signal correlations present in the training data. Capturing the full spectrum of noise distributions in real world images to prevent overfitting is impractical and even unrealistic.
@@
  Refer to caption PSNR: 32.73 dB Refer to caption (b) Gaussian Added   Refer to caption PSNR: 34.61 dB Refer to caption (c) Translated (Ours)
Figure 1: Denoising results of Gaussian-trained Restormer [36], applied to (a) original real noisy image, (b) noisy image with added Gaussian noise. We can observe that simply adding Gaussian noise to a real noisy image improves the denoising performance. (c) showcases our framework which translates real noise into Gaussian noise, enabling the denoising network to achieve even higher quality results. Zoom in for better details.
@@
  • We propose a novel noise translation network for robust image denoising, which converts unknown complex noise in an input image into a Gaussian noise. Our lightweight noise translation network can be applied to any Gaussian-pretrained denoising networks.
  • We employ well-motivated loss functions and architecture for the noise translation network. The proposed approach guides the noise distribution of the input image to Gaussian distribution both implicitly and explicitly by rigorously leveraging the mathematical properties.
@@ 2 Related Works
In recent years, deep learning has led to significant progress in image denoising, achieving impressive results by leveraging paired noisy and clean images for training. DnCNN [39] pioneered the use of CNNs for image denoising, which paved the way for further advancements involving residual learning [8, 19, 41], attention mechanisms [18, 41], and transformer models [36, 38]. Despite its success, acquiring the noisy-clean pairs required for supervised training remains a significant challenge. To address this, self-supervised approaches [16, 14, 2, 27, 17] have emerged to train networks using only noisy images, but these models typically perform considerably worse than their supervised counterparts. Additionally, zero-shot approaches [28, 10, 22] have been proposed for image denoising even without training dataset, but they require substantial computational cost at inference, making them impractical for real-time applications. In contrast to these methods that aim to reduce dependency on supervised data, our approach leverages supervised data but focuses on achieving good generalization performance with limited data.
@@
Generalization is a critical challenge in image denoising, as the performance of denoising models often degrades when encountering noise characteristics that were not seen during training. To handle unseen noise type or levels, DnCNN [39] employed blind Gaussian training to adapt to various noise levels, while Mohan et al. [25] designed a bias-free network to prevent overfitting to noise levels in the training set. More recent works employed masking-based learning [5] or leveraged the pre-trained CLIP encoder [7] to prevent overfitting by encouraging the model to understand global context rather than relying on local patterns. While these approaches enhance robustness to unseen noise, they often struggle to produce high-quality image restoration, particularly in complex real-world scenarios.
To address real-world noise, researchers have focused on constructing training datasets that closely resemble real noise distribution. This includes collecting real clean-noisy image pairs [1, 31, 33] and learning to generate realistic noise through data augmentation [34, 11, 4] or adversarial attacks [32, 29]. However, these approaches are limited to the noise distributions represented in the training dataset and fail to generalize effectively to unseen OOD noise. Recently, LAN [12] was introduced to address unseen noisy images by applying a learnable pixel-wise offset optimized at test time. However, this approach requires adjusting pixel-level parameters during inference, which limits scalability as it significantly increases memory and time demands when the input image size is large. In contrast, our framework does not require test-time training, allowing for scalable and efficient denoising at test time.
3 Method
@@
In this section, we present a robust image denoising framework featuring a novel noise translation process, designed to effectively handle diverse unseen noise. Our framework consists of: (1) a noise translation network to transform the arbitrary noise in an input image into ideal Gaussian noise, and (2) a denoising network to remove the translated noise to produce a clean output.
3.1 Image Denoising Network
First of all, we train a denoising network. Our image denoising network aims to recover a clean image from a noisy input, which can be mathematically formulated as
$\hat{I}=\mathcal{D}({I};\boldsymbol{\theta}),$ (1)
where $\mathcal{D}(\cdot;\boldsymbol{\theta})$ denotes an image denoising network parameterized by $\boldsymbol{\theta}$, and $I,\hat{I}\in\mathbb{R}^{H\times W\times C}$ represent a noisy input image and its corresponding denoised output, respectively.
@@ 3.1 Image Denoising Network
The goal of supervised training for our denoising network is to ensure that the denoised output $\hat{I}$, to closely match the ground-truth clean image $I_{\text{GT}}$. To achieve this, the model parameters $\boldsymbol{\theta}$ are optimized by minimizing the following loss function:
$\mathcal{L}=\|\mathcal{D}(I;\boldsymbol{\theta})-{I_{\text{GT}}}\|.$ (2)
Our method is model agnostic, allowing us to use existing denoising models and focus only on how to effectively remove Gaussian noise. To train our image denoising network to eliminate Gaussian noise, we utilize a training dataset consisting of clean images paired with their corrupted versions with synthetic additive Gaussian noise. We additionally use Gaussian-augmented real noisy-clean image pairs, where the noisy images are further corrupted with Gaussian noise. In our framework, image denoising network is optimized with the loss function in (2) to make it specialized in removing Gaussian noise.
@@ 3.2 Noise Translation Network
Figure 2 illustrates the overall training pipeline of the noise translation network. Formally, our framework first transforms a noisy image $I$ into an image with Gaussian noise $I_{\mathcal{T}}$, which is then fed into the denoising network to produce the final denoised output $\hat{I}_{\mathcal{T}}$, represented by
$\hat{I}_{\mathcal{T}}=\mathcal{D}(I_{\mathcal{T}};\boldsymbol{\theta}^{*})=\mathcal{D}(\mathcal{T}(I;\boldsymbol{\phi});\boldsymbol{\theta}^{*}),$ (3)
where $\mathcal{T}(\cdot,\boldsymbol{\phi})$ denotes the noise translation network with parameters $\boldsymbol{\phi}$, and $\mathcal{D}(\cdot,\boldsymbol{\theta}^{*})$ indicates the pretrained denoising network with parameters $\boldsymbol{\theta}^{*}$. Note that $\mathcal{D}(\cdot,\boldsymbol{\theta}^{*})$ is specialized in handling Gaussian noise, and its parameters are fixed during training the noise translation network. We next discuss how to train the noise translation network $\mathcal{T}(\cdot,\boldsymbol{\phi})$ by providing the following two loss terms.
@@ 3.2.1 Implicit Noise Translation Loss
Our goal is to transform an arbitrary noisy input image $I$ into a noise-translated image $I_{\mathcal{T}}$ that is well-suited for the pretrained denoising network, which is specialized for handling Gaussian noise. To achieve this, we optimize the noise translation network using a loss function, referred to as the implicit noise translation loss, which is designed to minimize the difference between the denoised image and the ground-truth clean image:
$\mathcal{L}_{\text{implicit}}=\|\hat{I}_{\mathcal{T}}-I_{\text{GT}}\|_{1}=\|\mathcal{D}(\mathcal{T}(I;\boldsymbol{\phi});\boldsymbol{\theta}^{*})-I_{\text{GT}}\|_{1},$ (4)
which guides the network to translate unseen noise into a form that the pretrained denoiser can handle effectively. A straightforward approach to train a noise translation network is to use real noisy-clean image pairs as $I$ and $I_{\text{GT}}$. To handle various noise levels in real-world scenarios that are lacking in the limited training set, we apply data augmentation by adding a random level of synthetic Gaussian noise to the noisy image $I$. This helps the noise translation network generalize more effectively across diverse noise conditions.
@@ 3.2.2 Explicit Noise Translation Loss
The implicit noise translation loss helps to translate the image into a form preferable for the pretrained denoising network, but it does not ensure that the noise distribution is transformed into an ideal Gaussian distribution, as it lacks direct control over the noise characteristics. To address this, we introduce an explicit loss function that directly guides the noise to follow Gaussian distribution.
Let $n_{\mathcal{T}}=I_{\mathcal{T}}-I_{GT}\in\mathbb{R}^{H\times W\times C}$ represent the translated noise, and let $n_{\mathcal{G}}\in\mathbb{R}^{H\times W\times C}$ be a random variable following a Gaussian distribution $\mathcal{N}(\hat{\mu},\hat{\sigma})$, where $(\hat{\mu},\hat{\sigma})$ denote the empirical mean and standard deviation calculated from all elements of $n_{\mathcal{T}}$. Our objective is to adjust the distribution of $n_{\mathcal{T}}$ to closely align with the distribution of $n_{\mathcal{G}}$. To achieve this, we utilize the Wasserstein distance to measure the difference between their distributions and employ it as a loss function to minimize:
@@
where $d_{W_{1}}(\cdot,\cdot)$ is 1-Wasserstein distance, also known as the Earth Mover’s Distance. To calculate this, we first flatten each channel in $n_{\mathcal{T}}$ and $n_{\mathcal{G}}$ over the spatial dimensions into one-dimensional vectors, and then sort them in an ascending order. Let $\mathbf{X}^{c}\equiv(X_{1}^{c},X_{2}^{c},\dots,X_{H\times W}^{c})$ and $\mathbf{Y}^{c}\equiv(Y_{1}^{c},Y_{2}^{c},\dots,Y_{H\times W}^{c})$ denote the ordered values of $n_{\mathcal{T}}$ and $n_{\mathcal{G}}$ for the $c^{\text{th}}$ channel, respectively. The 1-Wasserstein distance is then calculated by the following simple function of the order statistics [^1] Please refer to Appendix for the detailed proof.:
@@
$p_{R}(|F_{\mathcal{G}}^{c}(u,v)|;\sigma)=\frac{|F_{\mathcal{G}}^{c}(u,v)|}{\sigma^{2}}\exp\left(-\frac{|F_{\mathcal{G}}^{c}(u,v)|^{2}}{2\sigma^{2}}\right),$ (9)
which implies that $|F_{\mathcal{T}}^{c}(u,v)|$ should also follow a Rayleigh distribution to ensure that $n_{\mathcal{T}}$ is spatially uncorrelated. To this end, similar to Eqs. (5) and (6), we minimize the difference between the distributions of $|F_{\mathcal{T}}^{c}(u,v)|$ and $|F_{\mathcal{G}}^{c}(u,v)|$ by utilizing 1-Wasserstein distance, which is defined as
@@
where $\tilde{\mathbf{X}}^{c}\equiv(\tilde{X}^{c}_{1},\tilde{X}^{c}_{2},\dots,\tilde{X}^{c}_{H\times W})$ and $\tilde{\mathbf{Y}}^{c}\equiv(\tilde{Y}^{c}_{1},\tilde{Y}^{c}_{2},\dots,\tilde{Y}^{c}_{H\times W})$ are the sorted values of flattened magnitude of Fourier coefficients $|F_{\mathcal{T}}^{c}(u,v)|$ and $|F_{\mathcal{G}}^{c}(u,v)|$, respectively.
The full explicit noise translation loss is defined by $\mathcal{L}_{\text{spatial}}$ and $\mathcal{L}_{\text{freq}}$ as
@@
where $\beta$ is a hyperparameter that balances the contribution of the two Wasserstein distances. This loss function explicitly guides the translated noise to follow Gaussian distribution.
The total loss function for training the noise translation network is given by
@@
where $\alpha$ is a hyperparameter to control the influence of implicit and explicit loss terms.

### v2

@@ Abstract
Image denoising techniques based on deep learning often struggle with poor generalization performance to out-of-distribution real-world noise. To tackle this challenge, we propose a novel noise translation framework that performs denoising on an image with translated noise rather than directly denoising an original noisy image. Specifically, our approach translates unknown real-world noise into Gaussian noise, which is spatially uncorrelated and independent of image content, through our noise translation network. The translated noisy images are then processed by an image denoising network pretrained to effectively remove Gaussian noise, enabling robust and consistent denoising performance. We also design well-motivated loss functions and architectures for the noise translation network by leveraging the mathematical properties of Gaussian noise. Experimental results show that the proposed approach substantially improves robustness and generalizability, outperforming state-of-the-art methods across diverse benchmarks.
@@ 1 Introduction
Image denoising aims to restore the pure signal from noisy images and serves as a critical preprocessing step to improve the visual quality of input images, extending the applicability to various downstream tasks. Recent advances in deep learning facilitate significant performance improvement of image denoising models [46, 47, 10, 42, 44, 43, 6, 45, 36]. A common assumption in early approaches was that camera noise could be modeled as a Gaussian distribution [28, 46, 47], which simplified the process of generating noisy-clean image pairs by adding synthetic Gaussian noise. This allowed for the construction of large datasets that could be used to train denoising models in a supervised manner, playing a crucial role in advancing image denoising models.
While denoising models trained on synthetic datasets perform well in controlled environments, they struggle to generalize to real-world scenarios due to the fundamental differences between synthetic and real noise distributions [10]. In response, researchers collected clean-noisy image pairs from real images [1, 38, 40] to address realistic noises, but denoising models still tend to overfit to the specific noise-signal correlations present in the training data. Capturing the full spectrum of real-world noise distributions to prevent overfitting is impractical and even unrealistic.
@@
PSNR: 32.73 dB
(b) Gaussian Added
PSNR: 34.63 dB
(c) Translated (Ours)
Figure 2: Denoising results of Gaussian-trained Restormer [43]. From left to right: (1) real noisy image and its denoised output, (2) noisy image with added Gaussian noise and its denoised output, (3) noise-translated image with our framework and its corresponding denoised output, achieving superior quality.
@@
  • We propose a novel noise translation network for robust image denoising that converts unseen complex noise in an input image into Gaussian noise. Our lightweight noise translation network can seamlessly integrate with any Gaussian-pretrained denoising networks.
  • We employ well-motivated loss functions and architecture for the noise translation network. Our translation network transforms the noise distribution of the input image into a Gaussian distribution both implicitly and explicitly by leveraging key mathematical properties.
@@ 2 Related Works
In recent years, deep learning has led to significant progress in image restoration [6, 43, 13, 7, 35], achieving impressive results by leveraging paired noisy and clean images for training. DnCNN [46] pioneered the use of CNNs for image denoising, which paved the way for further advancements involving residual learning [9, 23, 48], attention mechanisms [22, 48], and transformer models [43, 45]. Despite its success, acquiring the noisy-clean pairs required for supervised training remains a significant challenge. To address this, self-supervised approaches [20, 18, 2, 32, 21] have emerged to train networks using only noisy images, but these models typically perform considerably worse than their supervised counterparts. Additionally, zero-shot approaches [33, 11, 27] have been proposed for image denoising even without a training dataset, but they require substantial computational cost at inference, making them impractical for real-time applications. While previous methods either rely heavily on supervised datasets or bypass them with performance trade-offs, we show that robust generalization is achievable even under limited supervision. This makes it suitable for real-world deployment, where collecting large-scale clean references remains a major bottleneck.
@@
Generalization is a critical challenge in image denoising, as the performance of denoising models often degrades when encountering noise characteristics that were not seen during training. To handle unseen noise type or levels, DnCNN [46] employed blind Gaussian training to adapt to various noise levels, while Mohan et al. [30] designed a bias-free network to prevent overfitting to noise levels in the training set. Recent works employed masking-based learning [5] or leveraged the pretrained CLIP encoder [8] to prevent overfitting by encouraging the model to understand global context rather than relying on local patterns. In addition, IDF [15] introduced an iterative dynamic filtering to better generalize to unseen noise. While these approaches enhance robustness to unseen noise, they often struggle to produce high-quality images, particularly in complex real-world scenarios.
To handle real-world noise, prior work has focused on building training datasets that reflect realistic noise distributions. This includes collecting real clean-noisy image pairs [1, 38, 40], simulating realistic noise via data augmentation [41, 12, 4], or leveraging adversarial training [39, 34]. However, these approaches are inherently limited by the noise distributions seen during training and often fail to generalize to out-of-distribution (OOD) noise. Alternatively, noise transformation methods such as the Anscombe transform [26] and pixel-shuffle down-sampling [50] aim to simplify denoising by converting complex noise into Gaussian-like noise. However, their reliance on fixed transformations limits their adaptability to diverse noise types. More recently, LAN [14] addressed unseen noise through test-time optimization of pixel-wise offsets. While effective, its per-pixel updates hinder scalability to images larger than $\times256$. In contrast, our approach removes the need for test-time training and enables efficient, scalable denoising under a wide range of noise conditions.
3 Image Denoising through Noise Translation
@@
This section presents our robust image denoising framework with a novel noise translation process, designed to handle diverse real-world noise. Our framework first transforms a noisy image $I$ into an image with Gaussian noise $I_{\mathcal{T}}$, which is then fed into the denoising network to produce the final denoised output $\hat{I}_{\mathcal{T}}$. The whole process is given by
$\hat{I}_{\mathcal{T}}=\mathcal{D}(I_{\mathcal{T}};\boldsymbol{\theta})=\mathcal{D}(\mathcal{T}(I;\boldsymbol{\phi});\boldsymbol{\theta}),$ (1)
where $\mathcal{T}(\cdot\,;\boldsymbol{\phi})$ denotes the noise translation network parametrized by $\boldsymbol{\phi}$, and $\mathcal{D}(\cdot\,;\boldsymbol{\theta})$ indicates the denoising network with parameters $\boldsymbol{\theta}$. Note that $\mathcal{D}(\cdot\,;\boldsymbol{\theta})$ is specialized in handling Gaussian noise in our framework. We next discuss how to train the denoising network $\mathcal{D}(\cdot\,;{\boldsymbol{\theta}})$ and noise translation network $\mathcal{T}(\cdot\,;\boldsymbol{\phi})$.
3.1 Image Denoising Network for Gaussian Noise
@@
We first train a denoising network by the standard supervised learning. Our denoising network, $\mathcal{D}(\cdot\,;\boldsymbol{\theta})$, aims to remove the Gaussian noise in an input image $I$ and reconstruct a clean image close to the ground-truth $I_{\text{GT}}$. To achieve this, the model parameters $\boldsymbol{\theta}$ are optimized by minimizing the following loss function:
$\mathcal{L}=\|\mathcal{D}(I;\boldsymbol{\theta})-{I_{\text{GT}}}\|_{1}.$ (2)
For training, the noisy input image $I$ is corrupted by synthetic Gaussian noise. Detailed procedures for dataset construction are provided in Section 4.1.
@@ 3.2 Noise Translation Network
We now describe the training procedure for the noise translation network $\mathcal{T}(\cdot;\boldsymbol{\phi})$. During this phase, the parameters of the denoising network $\mathcal{D}(\cdot\,;\boldsymbol{\theta}^{*})$, trained for Gaussian noise removal, are kept fixed to provide a stable supervisory signal. The overall training pipeline is illustrated in Figure 3.
@@ 3.2.1 Implicit Noise Translation Loss
The objective is to transform an arbitrary noisy input image $I$ into a noise-translated image $I_{\mathcal{T}}$ that is well-suited for the pretrained denoising network specialized in handling Gaussian noise. To this end, we optimize the noise translation network using the implicit noise translation loss, which minimizes the discrepancy between the denoised output and the corresponding clean ground-truth image:
$\mathcal{L}_{\text{implicit}}=\|\hat{I}_{\mathcal{T}}-I_{\text{GT}}\|_{1}=\|\mathcal{D}(\mathcal{T}(I;\boldsymbol{\phi});\boldsymbol{\theta}^{*})-I_{\text{GT}}\|_{1}.$ (3)
This loss function encourages the translation network to map arbitrary unseen noise into a form that can be effectively handled by the pretrained Gaussian Denoiser $\mathcal{D}(\cdot\,;\boldsymbol{\theta}^{*})$. A natural choice for constructing training pairs $(I,I_{\text{GT}})$ for training the noise translation network is to use real noisy-clean image pairs. Accordingly, we exclusively use noisy-clean pairs from the real noise dataset SIDD [1].
@@ 3.2.2 Explicit Noise Translation Loss
Although the implicit noise translation loss encourages the translated image to be well-aligned with the pretrained denoising network, it does not explicitly enforce that the translated noise follows an ideal Gaussian distribution, as it lacks direct control over the noise characteristics. To address this, we introduce an explicit loss that promotes the translated noise to follow a Gaussian distribution.
Let $n_{\mathcal{T}}=I_{\mathcal{T}}-I_{GT}\in\mathbb{R}^{H\times W\times C}$ represent the translated noise, and let $n_{\mathcal{G}}\in\mathbb{R}^{H\times W\times C}$ be a random variable following a Gaussian distribution $\mathcal{N}(\hat{\mu},\hat{\sigma}^{2})$, where $\hat{\mu}$ and $\hat{\sigma}$ denote the empirical mean and standard deviation calculated from all elements of $n_{\mathcal{T}}$, respectively. We adjust the distribution of $n_{\mathcal{T}}$ to align with the distribution of $n_{\mathcal{G}}$. To this end, we employ the Wasserstein distance between the two distributions as a loss function, which is given by
@@
where $d_{W_{1}}(\cdot,\cdot)$ is 1-Wasserstein distance, also known as the Earth Mover’s Distance. To compute this, we first flatten each channel in $n_{\mathcal{T}}$ and $n_{\mathcal{G}}$ over the spatial dimensions into one-dimensional vectors, and then sort them in an ascending order. Let $\mathbf{X}^{c}\equiv(X_{1}^{c},X_{2}^{c},\dots,X_{H\times W}^{c})$ and $\mathbf{Y}^{c}\equiv(Y_{1}^{c},Y_{2}^{c},\dots,Y_{H\times W}^{c})$ denote the ordered values of $n_{\mathcal{T}}$ and $n_{\mathcal{G}}$ for the $c^{\text{th}}$ channel, respectively. The 1-Wasserstein distance is then given by the following simple function [^1] Please refer to Appendix for the detailed proof.:
@@
$\hskip-5.406ptp_{R}(|F_{\mathcal{G}}^{c}(u,v)|;\sigma)=\frac{|F_{\mathcal{G}}^{c}(u,v)|}{\sigma^{2}}\exp\left(-\frac{|F_{\mathcal{G}}^{c}(u,v)|^{2}}{2\sigma^{2}}\right),$ (8)
which implies that $|F_{\mathcal{T}}^{c}(u,v)|$ should also follow a Rayleigh distribution, guiding $n_{\mathcal{T}}$ to be spatially uncorrelated. To this end, similar to Eqs. (4) and (5), we minimize the difference between the distributions of $|F_{\mathcal{T}}^{c}(u,v)|$ and $|F_{\mathcal{G}}^{c}(u,v)|$ by utilizing 1-Wasserstein distance, which is defined as
@@
where $\tilde{\mathbf{X}}^{c}\equiv(\tilde{X}^{c}_{1},\tilde{X}^{c}_{2},\dots,\tilde{X}^{c}_{H\times W})$ and $\tilde{\mathbf{Y}}^{c}\equiv(\tilde{Y}^{c}_{1},\tilde{Y}^{c}_{2},\dots,\tilde{Y}^{c}_{H\times W})$ are the sorted values of flattened magnitude of Fourier coefficients $|F_{\mathcal{T}}^{c}(u,v)|$ and $|F_{\mathcal{G}}^{c}(u,v)|$, respectively. The full explicit noise translation loss is defined by $\mathcal{L}_{\text{spatial}}$ and $\mathcal{L}_{\text{freq}}$ as
@@
where $\beta$ controls the contribution of the two Wasserstein distances. This loss function explicitly guides the translated noise to follow Gaussian distribution. The total loss function for training the noise translation network is given by
@@
where $\alpha$ balances the influence of two loss terms.